i do not know what to do on this problem.
How much 40% antifreeze solution must be mixed with a 80% antifreeze solution to produce 20 gallons of a 70% disinfectant solution? Round answer to the nearest gallon.
? gallons 40% solution, ? gallons 80% solution

Answer by checkley77(12844)

.40x+.80(20-x)=.70*20
.40x+16-.80x=14
-.40x=14-16
-.40x=-2
x=-2/.40
x=5 gal of 40% antifreeze is used.
20-5=15 gal of 80% anti freeze is used.
[5] gallons 40% solution, [15] gallons 80% solution
Proof:
.4*5+.8*15=.7*20
2+12=14
14=14